Determine the speed with which the Earth would have to rotate on its axis so that a person on the equator would weigh(3/5)th as much as present.take the equatorial radius as 6400 km.

Grade:11

1 Answers

Arun
6 years ago
Dear Saurabh
 
 

True weight at the equator, W = mg

Observed weight at the equator,

W' = mg' = (3/5)mg

at the equator λ = 0

 

mg' = mg - mRw² cosλ 

3mg/5 = mg - mRw². Cos 0

 = mg - mRw²

 

mRW² = 2mg/5

 

w = 7.8 * 10^-4 rad/ sec

 

Regards

Arun(askIITians forum expert)